OpenStudy (anonymous):

what are the intercepts of y=-0.75x+3.0

OpenStudy (mathmate):

Welcome to Openstudy! The intercepts of a line are the points where the line intersects the x- and y-axes. At the x-intercept, y=0. Similarly at the y-intercept, x=0.

OpenStudy (mathmate):

@robinson312

OpenStudy (anonymous):

so wats the answere 0,0

OpenStudy (mathmate):

So all you need to do is to put x=0 in the equation and solve for y. That would the the y-intercept. Similarly for x.

OpenStudy (anonymous):

so do 0=-0.75*0+3.0

OpenStudy (mathmate):

Exactly! What do you get for the y-intercept?

OpenStudy (mathmate):

Oops, you can only put either x or y to zero at a time, so 0=-0.75*y+3.0 and solve for y.

OpenStudy (mathmate):

Oops again: 0=-0.75*x+3.0 and solve for x. After that, you put x=0 and solve for y: y=-0.75*0+3.0 That will give you the two intercepts.

OpenStudy (anonymous):

wat I got for x is 3

OpenStudy (mathmate):

0=-0.75x+3 0.75x=3 x=3/0.75=4 is what I got.

OpenStudy (anonymous):

that's what I got

OpenStudy (tkhunny):

y=-0.75x+3.0 My favorite y=-(3/4)x+3.0 (3/4)x + y = 3 \(\dfrac{x}{4} + \dfrac{y}{3} = 1\) In the "Intercept Form", the intercepts are kind of just staring at you in those denominators.

OpenStudy (anonymous):

ok whats the y intercept of y=7

OpenStudy (tkhunny):

Seriously? Where does y = 7 cross the y axis? Let's go with (0,7).

OpenStudy (anonymous):

ok whats the intercept of 6x+8y=24